2. A loan of $350,000 is to be paid back to the lender in equal yearly payments during a period of 5 years. The loan terms were made using an interest rate of 12% compounded quarterly. Determine the amount of each yearly payment that will be made by the borrower. 123
Answer #1

yearly payment = Y

time period = 5 years

yearly interest rate = (1 + 12%/4)4 - 1 = 12.55%

Y/(1 + 12.55%) + .... + Y/(1 + 12.55%)5 = 350000

Y = $ 98421.9